# LASER: A Neuro-Symbolic Framework for Learning Spatio-Temporal Scene Graphs with Weak Supervision

### v5

@@ Abstract
Supervised approaches for learning spatio-temporal scene graphs (STSG) from video are greatly hindered due to their reliance on STSG-annotated videos, which are labor-intensive to construct at scale. Is it feasible to instead use readily available video captions as weak supervision? To address this question, we propose LASER, a neuro-symbolic framework to enable training STSG generators using only video captions. LASER employs large language models to first extract logical specifications with rich spatio-temporal semantic information from video captions. LASER then trains the underlying STSG generator to align the predicted STSG with the specification. The alignment algorithm overcomes the challenges of weak supervision by leveraging a differentiable symbolic reasoner and using a combination of contrastive, temporal, and semantics losses. The overall approach efficiently trains low-level perception models to extract a fine-grained STSG that conforms to the video caption. In doing so, it enables a novel methodology for learning STSGs without tedious annotations. We evaluate our method on three video datasets: OpenPVSG, 20BN, and MUGEN. Our approach demonstrates substantial improvements over fully-supervised baselines. On OpenPVSG, LASER achieves a unary predicate prediction accuracy of $27.78\%~{}(+12.65\%)$ and a binary recall@5 of $0.42~{}(+0.22)$. Furthermore, LASER exceeds baselines by $7\%$ on 20BN and $5.2\%$ on MUGEN in terms of overall predicate prediction accuracy.
@@ 3.2 Spatio-Temporal Specification Language (STSL)
[…]
@@ 3.4 Spatio-Temporal Alignment Checking
$\begin{array}[]{rrrl}\text{(Formula)}&\varphi&::=&a(\overline{t})~{}|~{}\varphi_{1}\wedge\varphi_{2}~{}|~{}\varphi_{1}\vee\varphi_{2}~{}|~{}\neg\varphi\\ &&~{}|{}&\bigcirc\varphi~{}|~{}\varphi_{1}~{}\mathbf{U}~{}\varphi_{2}~{}|~{}\square\varphi~{}|~{}\lozenge\varphi\\ \text{(Specification)}&\psi&::=&\exists v_{1},\dots,v_{k},\text{s.t.~{}}\varphi\\ \end{array}$
@@
$\begin{array}[]{rrlcl}\langle w,s\rangle&\models&\psi&\text{iff}&\exists\Gamma,\langle\Gamma,w,s\rangle\models\varphi\\ \langle\Gamma,w,s\rangle&\models&a(\bar{t})&\text{iff}&a(\bar{c})\in w[s]~{}\wedge~{}\bar{c}=\text{subst}_{\Gamma}(\bar{t})\\ \langle\Gamma,w,s\rangle&\models&\varphi_{1}\wedge\varphi_{2}&\text{iff}&\langle\Gamma,w,s\rangle\models\varphi_{1}~{}\wedge\\ &&&&\langle\Gamma,w,s\rangle\models\varphi_{2}\\ \langle\Gamma,w,s\rangle&\models&\neg\varphi&\text{iff}&\langle\Gamma,w,s\rangle\not\models\varphi\\ \langle\Gamma,w,s\rangle&\models&\bigcirc\varphi&\text{iff}&\langle\Gamma,w,s+1\rangle\models\varphi\\ \langle\Gamma,w,s\rangle&\models&\varphi_{1}\mathbf{U}\varphi_{2}&\text{iff}&\exists i.s\leq i\land\langle\Gamma,w,i\rangle\models\varphi_{2}\\ &&&&\forall k.s\leq k<i,\langle\Gamma,w,k\rangle\models\varphi_{1}\\ \\ \end{array}$
Figure 5: Formal semantics of STSL. $\langle w,s\rangle\models\psi$ means the STSL specification $\psi$ is aligned with the ST-SG $w$ starting from time $s$. We use $w\models\psi$ as an abbreviation for $\langle w,1\rangle\models\psi$.
@@
$\Pr(\mathbf{r}\models\psi)=\textstyle\sum_{w\in\mathcal{P}(\mathbf{r})}\Pr(w)\cdot\mathbbm{1}[w\models\psi],\quad\Pr(w)=\textstyle\prod_{f\in w}\Pr(f)\textstyle\prod_{f^{\prime}\in\mathbf{r}\backslash w}(1-\Pr(f^{\prime}))$ (4)
@@ 4.1 OpenPVSG Dataset
is-bendableis-fluidis-holdableis-rigidis-tearableaboveattachedbehindbrokenclosecloseddeformedemptyfarfitsfoldedhas-holehighinin-frontleftlownext-toonon-surfaceopenrightstackedstretchedtorntwistedunderuprightvisible000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111F1 ScoreGPALASER-PLASER
@@ 4.3 MUGEN Dataset
Experimental Setup. In the MUGEN dataset, each video is paired with a natural language description that can be represented as an ordered list of actions $[a_{1},a_{2},\dots,a_{n}]$. The corresponding STSL specifications are naturally expressed in the logical form $a_{1}\mathbf{U}a_{2}\dots\mathbf{U}a_{n}$. LASER uses these generated specifications, derived from natural language captions via a crafted semantic parser, as weak supervision labels. We evaluate the model’s performance based on action prediction accuracy.
@@ B.5 MUGEN
Learning Setup We train the model on $5000$ training datapoints, and $12,851$ test datapoints. The contrastive loss is obtained over a batch size 3; the violation loss is constructed with regard to the axioms in GPA Migimatsu & Bohg (2022). We train the model with a learning rate of $0.0001$, violation weight $0.01$, number of epochs $100$, and batch size $3$. The Scallop reasoning engine setup is using difftopkproofs provenance with $k$ set to $5$. We construct the temporal specification from the given text description using a simple heuristic-based semantic parser. We first extract an ordered action list $[a_{1},a_{2},\dots,a_{n}]$ from the text with heuristics, then construct the temporal specification as $a_{1}\mathbf{U}a_{2}\dots\mathbf{U}a_{n}$. This specification means the actions are performed one by one with no intermediate gaps. The generated specifications are used as ground truth programmatic STSL labels. As the MUGEN video data is very different from the natural videos, we train a classification model from scratch. Our neural perception model consists of an S3D video encoding model pretrained on Kinetics 400 Kay et al. (2017), and MLP layers to classify the actions. The video is first passed through the S3D Xie et al. (2018) model for frame-based embeddings. Then all the embeddings that occur in a video clip are concatenated as the input to a $2$-layer MLP classifier. We set the number of frames per clip to $2$, and the batch size to $3$ due to the hardware limitation. The classifier thus produces clip-wise video representations. We employ contrastive loss and semantic loss for training. More details are shown in Appendix B.5.
@@
Quantitative Study. We evaluate our action prediction performance on MUGEN and compare it to caption supervised VT-TWINS Ko et al. (2022), TempCLR Yang et al. (2023b), and a directly supervised baseline (Supervised). We provide all baselines with extra annotations on the start- and end-frame of each clip-based textual description. As the ground truth actions for the video are different from the text description, we use heuristics to obtain the start and end frames. For a fair comparison, we fine-tune an S3D backbone pre-trained using Kinetics 400 (the same as LASER) along with the baseline models. This label is used for training both VT-TWINS and fully supervised models. As shown in Table 4, LASER has better action prediction accuracy despite receiving less supervision than the baselines. Moreover, our weakly supervised model even achieves better accuracy than the Supervised method on #Data $=5000$. We further evaluate our approach on a downstream video-specification retrieval task. Given 3 videos and 3 specifications, we want to match the correctly aligned pairs, retrieve specification given video and vice versa. We denote the two tasks as spec-retrieval and video-retrieval respectively. Specifically, we evaluate the accuracy with which we infer whether the correct specification has the highest alignment score, and vice versa. LASER outperforms an embedding-based baseline SDSC Hayes et al. (2022) on both tasks (Table 9). Our approach can even identify actions that persist for a very short period of time, such as kill an enemy.

### v7

@@ Abstract
Supervised approaches for learning spatio-temporal scene graphs (STSG) from video are greatly hindered due to their reliance on STSG-annotated videos, which are labor-intensive to construct at scale. Is it feasible to instead use readily available video captions as weak supervision? To address this question, we propose LASER, a neuro-symbolic framework to enable training STSG generators using only video captions. LASER employs large language models to first extract logical specifications with rich spatio-temporal semantic information from video captions. LASER then trains the underlying STSG generator to align the predicted STSG with the specification. The alignment algorithm overcomes the challenges of weak supervision by leveraging a differentiable symbolic reasoner and using a combination of contrastive, temporal, and semantics losses. The overall approach efficiently trains low-level perception models to extract a fine-grained STSG that conforms to the video caption. In doing so, it enables a novel methodology for learning STSGs without tedious annotations. We evaluate our method on three video datasets: OpenPVSG, 20BN, and MUGEN. Our approach demonstrates substantial improvements over fully-supervised baselines. On OpenPVSG, LASER achieves a unary predicate prediction accuracy of $27.78\%~(+12.65\%)$ and a binary recall@5 of 0.42(+0.22)0.42~(+0.22). Furthermore, LASER exceeds baselines by $7\%$ on 20BN and $5.2\%$ on MUGEN in terms of overall predicate prediction accuracy.
@@ 3.2 Spatio-Temporal Specification Language (STSL)
Note that an argument to the predicate drop is a wildcard (_), since we do not specify where does the box drops from. This formula might seem too strict since it requires the two events to be consecutive. To make the specification more natural, one can change the above formula to $\lozenge(\texttt{touch(h, b)}\wedge\lozenge\texttt{drop(b, \_)})$”. Here, the two events, touch and drop, need to happen in chronological order but are not required to be consecutive.
@@ 3.4 Spatio-Temporal Alignment Checking
[…]
Figure 5: Formal semantics of STSL. $\langle w,[s:t]\rangle\models\psi$ means the STSL specification $\psi$ is aligned with the ST-SG $w$ starting from time $s$ till time $t$. We use $w\models\psi$ as an abbreviation for $\langle w,[1:m]\rangle\models\psi$, where $m$ is the full video length.
@@
[…]
@@ 4.1 OpenPVSG Dataset
is-bendableis-fluidis-holdableis-rigidis-tearableaboveattachedbehindbrokenclosecloseddeformedemptyfarfitsfoldedhas-holehighinin-frontleftlownext-toonon-surfaceopenrightstackedstretchedtorntwistedunderuprightvisible00.20.20.40.40.60.60.80.811F1 ScoreGPALASER-PLASER
@@ 4.3 MUGEN Dataset
Experimental Setup. In the MUGEN dataset, each video is paired with a natural language description that can be represented as an ordered list of actions $a_{1},a_{2},\dots,a_{n}]$. The corresponding STSL specifications are naturally expressed in the logical form $a_{1}\mathbf{U}a_{2}\dots\mathbf{U}a_{n}$. LASER uses these generated specifications, derived from natural language captions via a crafted semantic parser, as weak supervision labels. We evaluate the model’s performance based on action prediction accuracy.
@@ B.5 MUGEN
Learning Setup We train the model on $5000$ training datapoints, and $12,851$ test datapoints. The contrastive loss is obtained over a batch size 3; the violation loss is constructed with regard to the axioms in GPA Migimatsu & Bohg (2022). We train the model with a learning rate of $0.0001$, violation weight $0.01$, number of epochs $100$, and batch size $3$. The Scallop reasoning engine setup is using difftopkproofs provenance with $k$ set to $5$. We construct the temporal specification from the given text description using a simple heuristic-based semantic parser. We first extract an ordered action list $a_{1},a_{2},\dots,a_{n}]$ from the text with heuristics, then construct the temporal specification as $a_{1}\mathbf{U}a_{2}\dots\mathbf{U}a_{n}$. This specification means the actions are performed one by one with no intermediate gaps. The generated specifications are used as ground truth programmatic STSL labels. As the MUGEN video data is very different from the natural videos, we train a classification model from scratch. Our neural perception model consists of an S3D video encoding model pretrained on Kinetics 400 Kay et al. (2017), and MLP layers to classify the actions. The video is first passed through the S3D Xie et al. (2018) model for frame-based embeddings. Then all the embeddings that occur in a video clip are concatenated as the input to a $2$-layer MLP classifier. We set the number of frames per clip to $2$, and the batch size to $3$ due to the hardware limitation. The classifier thus produces clip-wise video representations. We employ contrastive loss and semantic loss for training. More details are shown in Appendix B.5.
@@
Quantitative Study. We evaluate our action prediction performance on MUGEN and compare it to caption supervised VT-TWINS Ko et al. (2022), TempCLR Yang et al. (2023b), and a directly supervised baseline (Supervised). We provide all baselines with extra annotations on the start- and end-frame of each clip-based textual description. As the ground truth actions for the video are different from the text description, we use heuristics to obtain the start and end frames. For a fair comparison, we fine-tune an S3D backbone pre-trained using Kinetics 400 (the same as LASER) along with the baseline models. This label is used for training both VT-TWINS and fully supervised models. As shown in Table 4, LASER has better action prediction accuracy despite receiving less supervision than the baselines. Moreover, our weakly supervised model even achieves better accuracy than the Supervised method on #Data =5000=5000. We further evaluate our approach on a downstream video-specification retrieval task. Given 3 videos and 3 specifications, we want to match the correctly aligned pairs, retrieve specification given video and vice versa. We denote the two tasks as spec-retrieval and video-retrieval respectively. Specifically, we evaluate the accuracy with which we infer whether the correct specification has the highest alignment score, and vice versa. LASER outperforms an embedding-based baseline SDSC Hayes et al. (2022) on both tasks (Table 9). Our approach can even identify actions that persist for a very short period of time, such as kill an enemy.